OpenStudy (anonymous):

What is the equation of a parabola with a vertex (0,0) and a focus at (0,4)?

OpenStudy (lalaly):

Vertex Form for a parabola: (x-h)^2 = 4p(y-k) vertex at (0,0) means h = k = 0 You now have x^2 = 4py focus at (0,4) means the distance from vertex to focus = 4 p = 4 4p = 16 EQUATION: 16y = x^2 y = (1/16)x^2